# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

### v2

@@ V-A2 Domain Shift and Joint Training
[…]
Combined Training
[…]
Fine-Tuning
[…]
V-A3 Model Capacity and Transferability
It is generally established in the literature that pretraining on a large dataset can boost the performance of a model even in self-supervised settings [19]. However, in seismic applications, the benefits of pretraining are critically sensitive to the distributional similarity between the source and target domains—a phenomenon that reflects the strong coupling between geologic variability and model transferability. As shown in our experiments, models pretrained on FaultSeg3D or CRACKS and fine-tuned on Thebe underperform compared to models trained from scratch or initialized from ImageNet weights (Fig. 8). In such distributionally mismatched cases, the standard pretaining-finetuning strategy may hinder rather than help performance.
A factor that also plays a role in modulating generalization is model capacity. Fig. 8 shows that large models like segformer benefit from pretraining on CRACKS, while smaller models (e.g., hed, deeplab-m) generalize better when trained from scratch. The observation suggests that both data alignment and model capacity affect transfer effectiveness. This interaction between model size and pretraining benefits suggests that a higher model capacity or parameter count is not always better, but rather amplifies both the advantages and disadvantages of transfer. Large-capacity architectures can better exploit well-aligned pretrained features but are also more prone to overfitting to domain-specific artifacts when source and target distributions diverge. Smaller-capacity models, by contrast, may act as an implicit regularizer, limiting over-specialization and enabling better zero-shot transfer in mismatched scenarios, as can be observed in Fig. 8.
Furthermore, models respond differently to fine-tuning. When pretraining on FaultSeg3D and fine-tuning on Thebe, models degrade in CRACKS performance (Fig. 9, blue circle) shows that Thebe induces domain shifts that are hard to unlearn. Conversely, even though the finetuning dataset is distributionally closer to the target, Thebe-pretrained segformer and deeplab also degrade on CRACKS after finetuning on FaultSeg3D as shown in Fig.  9 (green circles). The asymmetric behavior highlights the difficulty of finding universally robust pretraining strategies.
[…]
TABLE II: Results on domain adaptation case study
Training configuration FDA DANN
Dice Hausdorff Dice Hausdorff
FaultSeg3D $\rightarrow$ Cracks 0.139 48.221 0.156 41.058
FaultSeg3D $\rightarrow$ Thebe 0.276 20.277 0.319 15.872
@@ V-B1 Window Size and Loss Function
We can see that two trends emerge in these experiments. First, when using Dice as a loss, models benefit consistently from larger window sizes, with performance generally improving as the spatial context grows. This is likely because larger patches provide the network with a more complete view of fault structures, enabling better continuity modeling across sections. Second, when using BCE loss, this trend does not hold: performance remains flat or slightly declines with larger windows. This discrepancy stems from the class imbalance inherent to fault delineation [140]. BCE tends to work best when foreground and background classes are more balanced, whereas Dice loss is explicitly designed to handle imbalance. Using smaller patches effectively “zooms in” on the sparse fault regions, increasing the proportion of fault pixels and improving BCE performance.
@@ V-C1 Sensitivity to Visual Structure
Distance-based metrics are generally more tolerant to the structure of the predicted fault. This behavior is illustrated in Fig.  11. Where Fig. 11(a) represents the ground truth, while 11(b) and 11(c) show the predictions of two different models. Although the prediction in Fig. 11(b) appears structurally closer to the ground truth, it receives significantly worse BCD and Hausdorff scores compared to the prediction in Fig. 11(c). Since distance-based metrics do not heavily weigh continuity, the inclusion of a few extra pixels around the fault can improve the BCD score even if those pixels lack proper structural alignment. Notably, these patterns are not anomalies, different models often generate such consistent outputs. Consequently, numerical metrics can be misleading and may display discrepancies between one another.
@@ V-C2 Fault Sparsity
Another issue with distance-based metrics is that they are designed to measure the quality of a single continuous object in the image, whereas faults can consist of multiple sparse objects. This makes distance-based metrics very sensitive to both the number of faults present in the ground truth and their sparsity. We showcase this issue in Fig.  12, where we consider two cases: one with a few faults and another with many faults. Noise is added to both images, and each is compared against its original version. The case with fewer faults contains many outliers compared to the case with many faults, since the added noise often lies far from any existing fault. As discussed in Section II-E, because BCD is bidirectional, it accounts for each added noise pixel by searching for its closest existing fault. Therefore, having fewer faults results in a much worse BCD score. DICE, while also penalizing noise, is more stable across different sparsity levels.
@@ V-C3 Contradictory Scores and Human Judgement
Although pixel-based metrics are more stable, they remain sensitive to slight pixel shifts. In Fig.  13, we show examples corresponding to two different models. The prediction in Fig. 13(c) looks visually closer to the ground truth, but receives a worse Dice score (0.1325 vs. 0.13421) and significantly better BCD and Hausdorff (40.782 vs. 115.436 and 29.0 vs 72.5). Since the faults in both predictions are poorly structured and spatially misaligned, the Dice coefficient (being overlap-based) penalizes them similarly. In contrast, distance-based metrics such as BCD and Hausdorff distances are more sensitive to the spatial coherence of the predicted faults, hence, they are more tolerant to the structure of the faults. These contradictions imply that some metrics may conflict with visual intuition or downstream utility, and point to the need for context-aware metric selection frameworks.
@@ VI Analysis Using Fault Characteristic Metrics
TABLE III: Summary statistics of the three datasets: CRACKS (A), Thebe (B), and FS (C). Values are reported as mean $\pm$ standard deviation.
@@
TABLE IV: Evaluation metrics when models are pretrained on Thebe and tested on Thebe, with or without finetuning on other datasets.
@@
When evaluating transfer setups where all models were pretrained on Thebe and subsequently finetuned on different datasets before being tested back on Thebe, we observe a clear correspondence between the statistical characteristics of the finetuning dataset and the error patterns in the predictions. Training and testing solely on Thebe produces the strongest alignment with the ground truth, as indicated by the highest strike similarity (0.907), reflecting that the model effectively captures the long and continuous fault structures characteristic of Thebe. However, this setup also leads to an overestimation of fault length (length ratio of 1.65) and moderate over-segmentation (segment ratio of 3.29), suggesting that the model tends to exaggerate continuity while artificially fragmenting longer structures.
@@
TABLE V: Key observations
@@
VII Conclusion
Beyond conventional metrics such as Dice or Hausdorff distance, our geometric and topological analysis demonstrated that models also absorb the structural biases of the datasets they are finetuned on: CRACKS-trained models tended to reproduce short, jagged, crossover-rich structures, while FaultSeg3D-trained models favored smoother but disconnected faults. Even when pretrained on the same source, prediction styles were shaped by the statistical properties of the finetuning dataset. These findings underscore that evaluation should not rely solely on pixel-level accuracy but must also account for structural plausibility and geological realism.

### v3

@@ V-A3 Model Capacity and Transferability
TABLE V: Results on domain adaptation case study
Training configuration FDA DANN Finetuning
Dice Hausdorff Dice Hausdorff Dice Hausdorff
FaultSeg3D $\rightarrow$ Cracks 0.139 48.221 0.156 41.058 0.375 5.577
FaultSeg3D $\rightarrow$ Thebe 0.276 20.277 0.319 15.872 0.158 51.649
@@ V-B1 Window Size and Loss Function
We can see that two trends emerge in these experiments. First, when using Dice as a loss, models benefit consistently from larger window sizes, with performance generally improving as the spatial context grows. This is likely because larger patches provide the network with a more complete view of fault structures, enabling better continuity modeling across sections. Second, when using BCE loss, this trend does not hold: performance remains flat or slightly declines with larger windows. This discrepancy stems from the class imbalance inherent to fault delineation [sudre2017generalised]. BCE tends to work best when foreground and background classes are more balanced, whereas Dice loss is explicitly designed to handle imbalance. Using smaller patches effectively “zooms in” on the sparse fault regions, increasing the proportion of fault pixels and improving BCE performance.
@@ V-C1 Sensitivity to Visual Structure
Distance-based metrics are generally more tolerant to the structure of the predicted fault. This behavior is illustrated in Fig.  12. Where Fig. 12(a) represents the ground truth, while 12(b) and 12(c) show the predictions of two different models. Although the prediction in Fig. 12(b) appears structurally closer to the ground truth, it receives significantly worse BCD and Hausdorff scores compared to the prediction in Fig. 12(c). Since distance-based metrics do not heavily weigh continuity, the inclusion of a few extra pixels around the fault can improve the BCD score even if those pixels lack proper structural alignment. Notably, these patterns are not anomalies, different models often generate such consistent outputs. Consequently, numerical metrics can be misleading and may display discrepancies between one another.
@@ V-C2 Fault Sparsity
Another issue with distance-based metrics is that they are designed to measure the quality of a single continuous object in the image, whereas faults can consist of multiple sparse objects. This makes distance-based metrics very sensitive to both the number of faults present in the ground truth and their sparsity. We showcase this issue in Fig.  13, where we consider two cases: one with a few faults and another with many faults. Noise is added to both images, and each is compared against its original version. The case with fewer faults contains many outliers compared to the case with many faults, since the added noise often lies far from any existing fault. As discussed in Section II-E, because BCD is bidirectional, it accounts for each added noise pixel by searching for its closest existing fault. Therefore, having fewer faults results in a much worse BCD score. DICE, while also penalizing noise, is more stable across different sparsity levels.
@@ V-C3 Contradictory Scores and Human Judgement
Although pixel-based metrics are more stable, they remain sensitive to slight pixel shifts. In Fig.  14, we show examples corresponding to two different models. The prediction in Fig. 14(c) looks visually closer to the ground truth, but receives a worse Dice score (0.1325 vs. 0.13421) and significantly better BCD and Hausdorff (40.782 vs. 115.436 and 29.0 vs 72.5). Since the faults in both predictions are poorly structured and spatially misaligned, the Dice coefficient (being overlap-based) penalizes them similarly. In contrast, distance-based metrics such as BCD and Hausdorff distances are more sensitive to the spatial coherence of the predicted faults, hence, they are more tolerant to the structure of the faults. These contradictions imply that some metrics may conflict with visual intuition or downstream utility, and point to the need for context-aware metric selection frameworks.
@@ VI Analysis Using Fault Characteristic Metrics
TABLE VI: Summary statistics of the three datasets: CRACKS (A), Thebe (B), and FS (C). Values are reported as mean $\pm$ standard deviation.
@@
TABLE VII: Evaluation metrics when models are pretrained on Thebe and tested on Thebe, with or without finetuning on other datasets.
@@
When evaluating transfer setups where all models were pretrained on Thebe and subsequently finetuned on different datasets before being tested back on Thebe, we observe a clear correspondence between the statistical characteristics of the finetuning dataset (Table VI) and the error patterns in the predictions (Table VII). Training and testing solely on Thebe produces the strongest alignment with the ground truth, as indicated by the highest strike similarity (0.907), reflecting that the model effectively captures the long and continuous fault structures characteristic of Thebe. However, this setup also leads to an overestimation of fault length (length ratio of 1.65) and moderate over-segmentation (segment ratio of 3.29), suggesting that the model tends to exaggerate continuity while artificially fragmenting longer structures.
@@
Motivated by these observations, we distill them into a self-contained, general “model pick list” to guide out-of-the-box use (Table IX): first, profile the target dataset by labeling its fault–network statistics as High/Medium/Low (H/M/L) relative to Table VI; then select the row whose triggers match that profile and adopt the corresponding pretrain $\rightarrow$ finetune setup; finally, confirm the choice against cross-setup behavior in Table VII and apply the suggested lightweight post-prediction fixes. This table is intentionally model-agnostic with respect to architecture and emphasizes geometry-aware selection aligning to fault length, segmentation, sinuosity, curvature, stepovers, and spacing so that practitioners can choose the most appropriate setup for their basin characteristics without additional tuning.
VII Conclusion
Beyond conventional metrics such as Dice or Hausdorff distance, our geometric and topological analysis demonstrated that models also absorb the structural biases of the datasets they are finetuned on: CRACKS-trained models tended to reproduce short, jagged, crossover-rich structures, while FaultSeg3D-trained models favored smoother but disconnected faults. Even when pretrained on the same source, prediction styles were shaped by the statistical properties of the finetuning dataset. These findings underscore that evaluation should not rely only on pixel-level accuracy but also account for structural plausibility and geological realism.
TABLE VIII: Key observations
@@ VII Conclusion
TABLE IX: Out-of-the-box model pick list based on dataset fault–network statistics (Table VI) and cross-setup behavior (Table VII). Label each metric as High/Medium/Low (H/M/L) relative to Table VI, then select the row that matches; confirm with Table VII.
Dataset profile (vs. Table VI) General pick rule (what the model should do) Setup Suggestion Simple rationale (from Table VII) Quick fixes after prediction
Tortuous & highly segmented with frequent stepovers
Triggers: Sinuosity H, Segments H, Stepover density H, Curvature L–M, Length M
[…]
Long, continuous master faults; sparse stepovers (few segments)
Triggers: Length H, Segments L, Stepover density L, Sinuosity L–M, Curvature M
[…]
Balanced / uncertain regime (no extremes) or new basin
Triggers: All metrics within $\pm 1\sigma$ of Thebe (Table VI) or mixed signals
Start with a neutral, orientation-faithful model (high strike similarity) before specializing. Thebe Only Highest strike similarity and stable curvature; safe default when unsure. Enforce minimum segment length; trim dead-ends; lightly straighten detours.
High curvature and high tortuosity (complex relays)
Triggers: Curvature H, Sinuosity H, Stepover density M, Segments M–H
[…]
Stepover-dominated networks with moderate tortuosity
Triggers: Stepover density H, Sinuosity M, Segments M–H, Length M, Curvature L–M
[…]
Abbreviations. H/M/L: high/medium/low relative to Table VI. Length ratio $L_{\text{pred}}/L_{\text{gt}}$; Sinuosity ratio $S_{\text{pred}}/S_{\text{gt}}$; Stepover ratio $D_{\text{stepover}}^{\text{pred}}/D_{\text{stepover}}^{\text{gt}}$; Strike similarity: cosine similarity of strike histograms (see Table VII). Use Table VI to assign H/M/L triggers; confirm with Table VII (ratios near 11, small differences).
@@ Acknowledgment
[Uncaptioned image] Jorge Quesada received his B.E. and M.S. degrees from the Pontifical Catholic University of Peru. He joined the Georgia Institute of Technology as a Machine Learning PhD student in the department of Electrical and Computer Engineering in 2021, where he is now part of the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). He is interested in leveraging machine learning and image processing tools to study the mechanisms underlying computer and biological vision.
[Uncaptioned image] Chen Zhou is a Ph.D. student in the School of Electrical and Computer Engineering at the Georgia Institute of Technology. He is currently a Graduate Research Assistant in the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). He is working in the fields of machine learning,and image and video processing. His research interests include trajectory prediction and learning from label disagreement for the applications of autonomous vehicle, and seismic interpretation.
[Uncaptioned image] Prithwijit Chowdhury received his B.Tech. degree from KIIT University, India, in 2020. He joined the Georgia Institute of Technology as an M.S. student in the School of Electrical and Computer Engineering in 2021 and is currently pursuing his Ph.D. as a researcher in The Center for Energy and Geo Processing (CeGP) and as a member of the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). His research interests lie in digital signal and image processing and machine learning with applications to geophysics. He is an IEEE Student Member and a published author, with several works presented at the IMAGE conference and published in the GEOPHYSICS journal.
[Uncaptioned image] Mohammad Alotaibi is a Ph.D. student in the School of Electrical and Computer Engineering at the Georgia Institute of Technology. He is a Graduate Research Assistant in the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). His research focuses on machine learning and image processing, with particular interest in domain adaptation for seismic and medical imaging applications.
[Uncaptioned image] Ahmad Mustafa received his PhD in Electrical and Computer Engineering at the Georgia Institute of Technology. His research interests span deep learning, computer vision, and signal processing for medical and geophysical applications. He served as a Staff Scientist at Occidental Petroleum, leading the research and deployment of large-scale computational models for susbsurface applications. His research has been published in leading, high impact journals. He was awarded the prestigious Roger P. Webb ECE GRA Excellence Award in recognition of his research contributions. He has been involved in various teaching and instructional activities in both academic and industrial settings. His teaching services at Georgia Tech earned him several accolades, such as the Outstanding Online Head TA Award, the ECE GTA Excellence Award and the ECE CREATION Award, respectively.
[Uncaptioned image] Yusufjon Kumakov is currently serving as an Assistant Lecturer at Tashkent State Technical University. He earned his Master’s degree in Petroleum Engineering from the Polytechnic University of Turin, Italy. His primary research interests lie in the application of machine learning techniques to geophysics, with a focus on seismic imaging and signal processing.
[Uncaptioned image] Mohit Prabhushankar received his Ph.D. degree in electrical engineering from the Georgia Institute of Technology (Georgia Tech), Atlanta, Georgia, 30332, USA, in 2021. He is currently a Postdoctoral Research Fellow in the School of Electrical and Computer Engineering at the Georgia Institute of Technology in the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). He is working in the fields of image processing, machine learning, active learning, healthcare, and robust and explainable AI. He is the recipient of the Best Paper award at ICIP 2019 and Top Viewed Special Session Paper Award at ICIP 2020. He is the recipient of the ECE Outstanding Graduate Teaching Award, the CSIP Research award, and of the Roger P Webb ECE Graduate Research Assistant Excellence award,all in 2022. He has delivered short courses and tutorials at IEEE IV’23, ICIP’23, BigData’23, WACV’24 and AAAI’24.
[Uncaptioned image] Ghassan AlRegib is currently the John and Marilu McCarty Chair Professor in the School of Electrical and Computer Engineeringat the Georgia Institute of Technology. In theOmni Lab for Intelligent Visual Engineering and Science (OLIVES), he and his groupwork on robustand interpretable machine learning algorithms, uncertainty and trust, and human in the loop algorithms. The group has demonstrated their work on a widerange of applications such as Autonomous Systems, Medical Imaging, and Subsurface Imaging. The group isinterested in advancing the fundamentals as well as the deployment of such systems in real-world scenarios. He has been issued several U.S.patents and invention disclosures. He is a Fellow of the IEEE. Prof. AlRegib is active in the IEEE. He served on the editorial board of several transactions and served as the TPC Chair for ICIP 2020, ICIP 2024, and GlobalSIP 2014. He was area editor for the IEEE Signal Processing Magazine. In 2008, he received the ECE Outstanding Junior Faculty Member Award. In 2017, he received the 2017 Denning Faculty Award for Global Engagement.He received the 2024 ECE Distinguished Faculty Achievement Award at Georgia Tech. He and his students received the Best Paper Award in ICIP 2019and the 2023 EURASIP Best Paper Award for Image communication Journal. In addition, one of their papers is the best paper runner-up at BigData 2024. In 2024, he co-founded the AI Makerspace at Georgia Tech, where any student and any community member can access and utilize AI regardless of their background.
[…]